The length of a rectangle is 3 times it's width. it's perimeter is 56cm. what is the area of the rectangle?
Answer by umnikmo(75)

P=2(L+W)
L=3W
P=2(W+3W)=8W=56 =>W=7
L=3W=21
S=WL=21*7=147